
At 22:00, CHIRON was rebooted because there was communication problems between Nightly observing script website and TCS.
At 23:00 , same problem.It was fixed quickly.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2023-05-14 | 8.0 | 0.0 | 0.25 | 2.0 | 0.0 | 10.25 |
Links
[1] http://www.ctio.noao.edu/noao/field-collection/field-nr-conditions/42500/delete?destination=print/node/20438
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-conditions/42500/edit?destination=print/node/20438
[3] mailto:thenry@astro.gsu.edu
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/42501/delete?destination=print/node/20438
[5] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/42501/edit?destination=print/node/20438
[6] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/20438?destination=print/node/20438
At the beginning of the night , the dome remained closed because it was cloudy. At 22:05 , the observing program started.
Objects : T0391903064, TOI-3378, CD-547010 and RKS1935-2259 were not observed.